# Unveiling the Role of Learning Rate Schedules via Functional Scaling Laws

### v1

@@ D.2 Proof for The Exponential-Decay LRS
Let $r=uT$. Then, by a change of variable, we obtain
$I_{2}=\frac{1}{4T^{1-1/\beta}}\int_{\frac{T}{M^{\beta}}}^{T}\frac{1-e^{-2r}-2re^{-2r}}{r^{1+1/\beta}}\mathop{}\!\mathrm{d}r=:\frac{1}{4T^{1-1/\beta}}\int_{\frac{T}{M^{\beta}}}^{T}q_{\beta}(r)\mathop{}\!\mathrm{d}r.$
It is easy to verify that for any $\beta\geqslant 1$, $\inf_{r\geqslant 0}q_{\beta}(r)\geqslant 0$ and $q_{\beta}(r)\eqsim r^{-1-1/\beta}$ when $r$ is sufficiently large. We refer to Figure 11 for an illustration of $q_{\beta}(\cdot)$.
Figure 11: Illustration of the function $q_{\beta}(\cdot)$.
  • When $T/M^{\beta}\leqslant 1$ and $T$ is sufficiently large such that $\int_{\frac{T}{M^{\beta}}}^{T}q_{\beta}(r)\mathop{}\!\mathrm{d}r\eqsim\beta$ and thus we have
    $I_{2}=\frac{\beta+o_{M,T}(1)}{4T^{1-1/\beta}}.$
  • When $T/M^{\beta}>1$, it holds for all $r\geqslant 1$ that $0.5\leqslant 1-e^{-2r}-2re^{-2r}\leqslant 1$ . Thus, there exists a $C_{T,M}\in[0.5,1]$ such that
    $I_{2}$ $=C_{T,M}\frac{1}{4T^{1-1/\beta}}\int_{\frac{T}{M^{\beta}}}^{T}r^{-1-1/\beta}\mathop{}\!\mathrm{d}r$
    $=\frac{C_{T,M}\beta}{4T^{1-1/\beta}}\left(\left(\frac{T}{M^{\beta}}\right)^{-1/\beta}-T^{-1/\beta}\right)=\frac{C_{T,M}\beta(M-1)}{4T}.$
Combining the two cases, we complete the proof. ∎
@@
$\varphi(\tau)=ae^{-\lambda\tau},\text{ with }\varphi(K)=b,$
@@
Given a large total compute budget $C\gg 1$, the optimal strategy for minimizing the final population risk—expressed in terms of the effective maximum learning rate $\gamma$, model size $M$, and data size DD—is given by:
@@ D.3 Proof for the WSD-Like LRS
$\int_{0}^{\infty}\mathcal{K}(T_{2}+t)\mathop{}\!\mathrm{d}t\lesssim\frac{\min\{M,T_{2}^{\frac{1}{\beta}}\}}{T_{2}}.$
@@
$\int_{0}^{\infty}\mathcal{K}(T_{2}+t)\mathop{}\!\mathrm{d}t$ $=\int_{0}^{\infty}\int_{M^{-\beta}}^{1}u^{1-\frac{1}{\beta}}e^{-2u(T_{2}+t)}\mathop{}\!\mathrm{d}u\mathop{}\!\mathrm{d}t$
$=\int_{M^{-\beta}}^{1}u^{1-\frac{1}{\beta}}e^{-2uT_{2}}\frac{\mathop{}\!\mathrm{d}u}{2u}$
$\eqsim\frac{1}{T_{2}^{1-\frac{1}{\beta}}}\int_{T_{2}M^{-\beta}}^{T_{2}}u^{-\frac{1}{\beta}}e^{-2u}\mathop{}\!\mathrm{d}u.$
Since the integral $\int_{0}^{\infty}u^{-\frac{1}{\beta}}e^{-2u}\mathop{}\!\mathrm{d}u$ is convergent, we have
$\int_{0}^{\infty}\mathcal{K}(T_{2}+t)\mathop{}\!\mathrm{d}t\lesssim\frac{1}{T_{2}^{1-\frac{1}{\beta}}}.$
When $T_{2}>M^{\beta}$, similarly we have
$\int_{0}^{\infty}\mathcal{K}(T_{2}+t)\mathop{}\!\mathrm{d}t$ $\eqsim M^{1-\beta}\int_{1}^{M^{\beta}}u^{-\frac{1}{\beta}}e^{-2u\frac{T_{2}}{M^{\beta}}}\mathop{}\!\mathrm{d}u.$
Let $p=\frac{T_{2}}{M^{\beta}}\geqslant 1$, we have
$\int_{0}^{\infty}\mathcal{K}(T_{2}+t)\mathop{}\!\mathrm{d}t$ $\eqsim\frac{M}{T_{2}}p\int_{1}^{M^{\beta}}u^{-\frac{1}{\beta}}e^{-2up}\mathop{}\!\mathrm{d}u$
$\lesssim\frac{M}{T_{2}}\int_{1}^{M^{\beta}}u^{-\frac{1}{\beta}}e^{-2u}\mathop{}\!\mathrm{d}u\eqsim\frac{M}{T_{2}}.$
Where the last line is because $pe^{-2up}$ is decreasing in $p$ when $u,p\geqslant 1$. ∎
@@
where $T=aK_{1}+(a-b)K_{2}/\log(a/b)$ is the total intrinsic training time, and $T_{2}=(a-b)K_{2}/\log(a/b)$ is the decay-phase intrinsic training time.
@@
$\int_{0}^{T(K)}\mathcal{K}(T(K)-t)\gamma_{\varphi}(t)\mathop{}\!\mathrm{d}t=\int_{0}^{T_{1}}\mathcal{K}(T(K)-t)a\mathop{}\!\mathrm{d}t+\int_{0}^{T_{2}}\mathcal{K}(T_{2}-t)(a-\lambda t)\mathop{}\!\mathrm{d}t,$
@@
$\int_{0}^{T_{2}}\mathcal{K}(T_{2}-t)(a-\lambda t)\mathop{}\!\mathrm{d}t\eqsim b+\frac{(a-b)\min\{M,T_{2}^{\frac{1}{\beta}}\}}{T_{2}}.$
@@
$\int_{0}^{T(K)}\mathcal{K}(T(K)-t)\gamma_{\varphi}(t)\mathop{}\!\mathrm{d}t$ $\eqsim\int_{0}^{T_{1}}\mathcal{K}(T(K)-t)a\mathop{}\!\mathrm{d}t+b+\frac{(a-b)\min\{M,T_{2}^{\frac{1}{\beta}}\}}{T_{2}}$
$\eqsim a\int_{0}^{T_{1}}\mathcal{K}(T_{2}+t)\mathop{}\!\mathrm{d}t+b+\frac{(a-b)\min\{M,T_{2}^{\frac{1}{\beta}}\}}{T_{2}}$
@@
    $\gamma_{\mathrm{opt}}\eqsim(C/\log C)^{-\frac{1+s\beta-\beta}{2+s\beta}},M_{\mathrm{opt}}\eqsim(C/\log C)^{\frac{1}{2+s\beta}},D_{\mathrm{opt}}\eqsim C^{\frac{1+s\beta}{2+s\beta}}(\log C)^{\frac{1}{2+s\beta}},(D_{1})_{\mathrm{opt}}\eqsim D,(D_{2})_{\mathrm{opt}}\eqsim D,$
@@ Appendix E Auxiliary Lemmas
For a semi-positive definite (SPD) matrix $\mathbf{A}$ and a random gaussian vector $\mathbf{x}\sim\mathcal{N}(0,\mathbf{H})$,
@@
Let $\mathbf{P}\preceq\mathbf{Q}$ be two positive semi-definite (p.s.d.) matrices. Then for any p.s.d. matrix $\mathbf{U}$, we have
@@
since $\mathbf{U}$ and $\mathbf{Q}-\mathbf{P}$ are both SPD matrices. ∎

### v3

@@ D.2 Proof for The Exponential-Decay LRS
Recall that
$I_{2}=\sum_{j=1}^{M}\!\left(-\frac{\lambda_{j}e^{-2\lambda_{j}T}}{2}+\frac{1-e^{-2\lambda_{j}T}}{4T}\right)=\frac{1}{4T}\sum_{j=1}^{M}\!\bigl(1-e^{-2\lambda_{j}T}-2\lambda_{j}Te^{-2\lambda_{j}T}\bigr),$
where $\lambda_{j}\eqsim j^{-\beta}$. Let $c_{1},c_{2}$ be constants such that $c_{1}j^{-\beta}\leqslant\lambda_{j}\leqslant c_{2}j^{-\beta}$.
Case 1. $T\geqslant M^{\beta}$.
In this regime, for every $j\leqslant M$ we have $\lambda_{j}T\geqslant c_{1}M^{-\beta}T\geqslant c_{1}$. Since $h(r)=1-e^{-2r}-2re^{-2r}$ is bounded on $c_{1},\infty)$, there exist constants $c_{3}<c_{4}$ such that $h(r)\in[c_{3},c_{4}]$ for all $r\geqslant c_{1}$. Thus
$\frac{c_{3}}{4}\frac{M}{T}\leqslant I_{2}\leqslant\frac{c_{4}}{4}\frac{M}{T}.$
Case 2. $T<M^{\beta}$.
Since $g(r)=1-e^{-2r}-2re^{-2r}\geqslant 2r^{2}e^{-2r}$ for $r>0$, we have
$I_{2}\geqslant\frac{1}{4T}\sum_{j=1}^{M}2\lambda_{j}^{2}T^{2}e^{-2\lambda_{j}T}=\frac{T}{2}\mathcal{K}_{M}(T)\eqsim\frac{1}{T^{1-\frac{1}{\beta}}},$
where the last equality follows from Lemma C.4 that $\mathcal{K}_{M}(t)\eqsim t^{-(2-1/\beta)}$ when $1\lesssim t\lesssim M^{\beta}$.
On the other hand, since the term $1-e^{-2\lambda_{j}T}-2\lambda_{j}Te^{-2\lambda_{j}T}$ is increasing in $\lambda_{j}$ and $\lambda_{j}\leqslant c_{2}j^{-\beta}$, we have
$I_{2}$ $\leqslant\frac{1}{4T}\sum_{j=1}^{M}\left(1-e^{-2c_{2}j^{-\beta}T}-2c_{2}j^{-\beta}e^{-2c_{2}j^{-\beta}T}\right)$
$\leqslant\frac{1}{4T}\int_{0}^{\infty}\left(1-e^{-2c_{2}x^{-\beta}T}-2c_{2}x^{-\beta}Te^{-2c_{2}x^{-\beta}T}\right)\mathop{}\!\mathrm{d}x$
${\tiny(u=2c_{2}Tx^{-\beta})\quad}$ $=\frac{(2c_{2})^{\frac{1}{\beta}}}{4T^{1-\frac{1}{\beta}}}\int_{0}^{\infty}\frac{1-e^{-u}-ue^{-u}}{u^{1+\frac{1}{\beta}}}\mathop{}\!\mathrm{d}u$
$\lesssim\frac{1}{T^{1-\frac{1}{\beta}}}.$
where the last inequality follows from the integral
$\int_{0}^{\infty}\frac{1-e^{-u}-ue^{-u}}{u^{1+\frac{1}{\beta}}}\mathop{}\!\mathrm{d}u<\infty,$
which is because the function $q(u):=\frac{1-e^{-u}-ue^{-u}}{u^{1+\frac{1}{\beta}}}$ satisfies $q(u)\sim u^{1-\frac{1}{\beta}}$ as $u\to 0$, and $q(u)\sim u^{-(1+\frac{1}{\beta})}$ as $u\to\infty$.
Combining the two cases, we obtain that
$I_{2}\eqsim\frac{\min(M,\,T^{1/\beta})}{T}.$
@@
$\varphi(\tau)=ae^{-\zeta\tau},\text{ with }\varphi(K)=b,$
@@
Given a large total compute budget $C\gg 1$, the optimal strategy for minimizing the final population risk – expressed in terms of the effective maximum learning rate $\gamma$, model size $M$, and data size $D$ – is given by:
@@ D.3 Proof for the WSD-Like LRS
$\int_{0}^{\infty}\mathcal{K}_{M}(T_{2}+t)\mathop{}\!\mathrm{d}t\lesssim\frac{\min\{M,T_{2}^{\frac{1}{\beta}}\}}{T_{2}}.$
@@
$\int_{0}^{\infty}\mathcal{K}_{M}(T_{2}+t)\mathop{}\!\mathrm{d}t$ $=\int_{0}^{\infty}\sum_{j=1}^{M}\lambda_{j}^{2}e^{-2\lambda_{j}(T_{2}+t)}\mathop{}\!\mathrm{d}t$
$=\sum_{j=1}^{M}\frac{\lambda_{j}}{2}e^{-2\lambda_{j}T_{2}}$
$\lesssim\frac{1}{T_{2}^{1-\frac{1}{\beta}}},$
where the last line can be shown using the same approach as in the proof of Lemma C.4.
On the other hand, we have
$\int_{0}^{\infty}\mathcal{K}_{M}(T_{2}+t)\mathop{}\!\mathrm{d}t$ $=\int_{0}^{\infty}\sum_{j=1}^{M}\lambda_{j}^{2}e^{-2\lambda_{j}(T_{2}+t)}\mathop{}\!\mathrm{d}t$
$=\sum_{j=1}^{M}\frac{\lambda_{j}}{2}e^{-2\lambda_{j}T_{2}}$
$=\frac{1}{2T_{2}}\sum_{j=1}^{M}\lambda_{j}T_{2}e^{-2\lambda_{j}T_{2}}\lesssim\frac{M}{T_{2}}\sup_{x>0}xe^{-2x}\lesssim\frac{M}{T_{2}},$
where the last line is because $\sup_{x>0}xe^{-2x}=\frac{1}{2e}$. ∎
@@
where $T=aK_{1}+(a-b)K_{2}/\log(a/b)$ is the total intrinsic training time, $T_{2}=(a-b)K_{2}/\log(a/b)$ is the decay-phase intrinsic training time, and we require that $T\gtrsim 1$, $T_{2}\gtrsim 1$.
@@
$\int_{0}^{T(K)}\mathcal{K}_{M}(T(K)-t)\gamma(t)\mathop{}\!\mathrm{d}t=\int_{0}^{T_{1}}\mathcal{K}_{M}(T(K)-t)a\mathop{}\!\mathrm{d}t+\int_{0}^{T_{2}}\mathcal{K}_{M}(T_{2}-t)(a-\zeta t)\mathop{}\!\mathrm{d}t,$
@@
$\int_{0}^{T_{2}}\mathcal{K}_{M}(T_{2}-t)(a-\zeta t)\mathop{}\!\mathrm{d}t\eqsim b+\frac{(a-b)\min\{M,T_{2}^{\frac{1}{\beta}}\}}{T_{2}}.$
@@
$\int_{0}^{T(K)}\mathcal{K}_{M}(T(K)-t)\gamma(t)\mathop{}\!\mathrm{d}t$ $\eqsim\int_{0}^{T_{1}}\mathcal{K}_{M}(T(K)-t)a\mathop{}\!\mathrm{d}t+b+\frac{(a-b)\min\{M,T_{2}^{\frac{1}{\beta}}\}}{T_{2}}$
$\eqsim a\int_{0}^{T_{1}}\mathcal{K}_{M}(T_{2}+t)\mathop{}\!\mathrm{d}t+b+\frac{(a-b)\min\{M,T_{2}^{\frac{1}{\beta}}\}}{T_{2}}$
@@
    $\gamma_{\mathrm{opt}}\eqsim(C/\log C)^{-\frac{1+s\beta-\beta}{2+s\beta}},M_{\mathrm{opt}}\eqsim(C/\log C)^{\frac{1}{2+s\beta}},D_{\mathrm{opt}}\eqsim C^{\frac{1+s\beta}{2+s\beta}}(\log C)^{\frac{1}{2+s\beta}},$
    $(D_{1})_{\mathrm{opt}}\eqsim D,(D_{2})_{\mathrm{opt}}\eqsim D,$
@@ Appendix E Auxiliary Lemmas
For any PSD matrix $\mathbf{A}$ and a random gaussian vector $\mathbf{x}\sim\mathcal{N}(0,\mathbf{H})$,
@@
Let $\mathbf{P}\preceq\mathbf{Q}$ be two PSD matrices. Then for any PSD matrix $\mathbf{U}$, we have
@@
since $\mathbf{U}$ and $\mathbf{Q}-\mathbf{P}$ are both PSD matrices. ∎